# From Spurious to Causal: Low-rank Orthogonal Subspace Intervention for Generalizable Face Forgery Detection

### v1

@@ Abstract
The generalization problem remains a critical challenge in face forgery detection. Some researches have discovered that “a backdoor path” in the representations from forgery-irrelevant information to labels induces biased learning, thereby hindering the generalization. In this paper, these forgery-irrelevant information are collectively termed spurious correlations factors. Previous methods predominantly focused on identifying concrete, specific spurious correlation and designing corresponding solutions to address them. However, spurious correlations arise from unobservable confounding factors, making it impractical to identify and address each one individually. To address this, we propose an intervention paradigm for representation space. Instead of tracking and blocking various instance-level spurious correlation one by one, we uniformly model them as a low-rank subspace and intervene in them. Specifically, we decompose spurious correlation features into a low-rank subspace via orthogonal low-rank projection, subsequently removing this subspace from the original representation and training its orthogonal complement to capture forgery-related features. This low-rank projection removal effectively eliminates spurious correlation factors, ensuring that classification decision is based on authentic forgery cues. With only 0.43M trainable parameters, our method achieves state-of-the-art performance across several benchmarks, demonstrating excellent robustness and generalization.
11footnotetext: Equal contribution.22footnotetext: Corresponding authors.
1 Introduction
@@
Figure 1: (a) An example of the impact of spurious correlations on Vanilla CLIP [38]. A backdoor path exists from forgery-irrelevant information (e.g. background style, identity information) to labels. (b) Causal intervention. We propose orthogonal low-rank projection to remove the spurious correlation subspace, disrupting the backdoor path.
@@
This paper explores this issue from a causal learning perspective. Visual representations contain both causal factors related to forgery and non-causal factors unrelated to forgery. When all parameters are frozen and only one trainable classification head is added, Vanilla CLIP [38] tends to treat non-causal factors (e.g., background, identity, and domain differences) as discriminative cues, thereby degrading detection performance (see Figure 1 (a.2)). From a causal perspective, these forgery-irrelevant factors easily form the backdoor path to labels, leading CLIP to shortcut learning and deviate from the correct causal path from forgery traces to labels (see Figure 1 (a.3)). A straightforward solution is to design suppression strategies for specific spurious correlation. However, spurious correlations are complex and variable, often arising from unobservable confounding factors. Such targeted approaches struggle to completely eliminate backdoor paths at the representation level.
To address this issue, we propose a novel method called SeLop (Spurious correlation elimination via Low-rank orthogonal projection for face forgery detection). SeLop performs a subspace-level unified intervention in CLIP’s visual representation space and constrains the decision evidence to align with forgery cues. The key idea of SeLop is to regard all spurious correlations factors as a compressible low-rank whole, as shown in Figure 1 (b). Within the same subspace, we perform holistic modeling and constraint on these factors and enforce orthogonal separation from forgery-related cues. This avoids per-case enumeration and local patching, and achieves overall removal of spurious correlations in the representation space. Specifically, we define a trainable basis matrix and perform QR decomposition on it to obtain the required low-rank orthogonal space. We then project the original CLIP visual space into the low-rank orthogonal space to obtain the spurious correlation subspace. Finally, we remove this subspace from the visual representation, allowing the true causal factors to be concentrated in the orthogonal complement space. Through end-to-end training with a very low parameter count of 0.43M, the model identifies spurious correlation factors and knows what should be removed in order to learn causal discriminant factors. The explicit orthogonality also maximizes the retention of CLIP’s rich pre-training knowledge while learning forgery-related features.
@@
    In this paper, we approach spurious correlations from a new perspective: rather than discovering a specific backdoor path, we consider the backdoor path relationships caused by spurious correlations as a whole and intervene accordingly.
@@
    We effectively intervene in CLIP’s original visual space through orthogonal low-rank removal with minimal parameters, enabling CLIP to learn to rely on the correct causal path for classification decisions.
@@
    Extensive experiments on several standard benchmarks demonstrate the superior generalization and robustness of our method. In particular, with only 0.43M trainable parameters, our method achieves state-of-the-art performance.
2 Related Work
2.1 Face Forgery Detection
@@
The rapid advancement of artificial intelligence technologies has spurred diversification in facial forgery techniques. Presently, improving generalizability of detection models remains a challenging issue in face forgery detection tasks. Previous work has focused on meticulously designing deep neural networks to learn “specific and fake patterns”, including biological signals (e.g., lip movement [13], gaze angle [36]), frequency anomalies (e.g., F2Trans [32], F3Net [37]), and identity disentanglement (e.g., CADDM [10]). However, due to the limited forgery patterns in training datasets, models often overfit and fail to generalize to unknown forgery techniques. To mitigate overfitting, [22, 41] proposed a pseudo-fake face generation strategy. Training models on custom datasets improved generalization to some extent. However, expecting models to learn all forgery patterns on custom data remains impractical.
@@
With the advent of large language models (LLMs), pre-trained visual foundation models (e.g., CLIP [38]) on massive data have demonstrated strong generalization across downstream tasks. Recent research has explored adapting CLIP for face forgery detection. UDD [11] designed a shuffle branch and a mix branch to break position and content biases of the CLIP. GM-DF [20] proposes a domain-aware hybrid expert modelling approach within a meta-learning framework, achieving generalized multi-scenario detection. RepDFD [24] injects perturbations into input images to reprogram pre-trained VLM models. Forensic-Adapter [6] introduces an adapter specifically to learn blending boundaries, achieving excellent generalization through interaction between CLIP and the adapter. Effort [45] employs Singular Value Decomposition (SVD) to decouple the feature space of attention, preserving pre-trained knowledge while learning forgery patterns.
2.2 Causal Learning
@@
In recent years, causal learning has attracted significant attention in the field of computer vision. Mitigating spurious correlations through causal intervention has proven a promising approach to improving cross-domain generalization [50]. Recently, some researchers have introduced causal learning into face forgery detection. UDD [11] identified position and content biases in CLIP [38], designing shuffling and mixing branches to break inherent biases and achieve great generalization.
@@
3 Problem Formulation
@@
In this paper, we formulate the problem of face forgery detection from a novel perspective of causal learning. Following [11], we utilize Structural Causal Model (SCM) [35] to define the relevant variables and factors, as shown in Figure 1. Let $U$ denote unobserved latent confounders, $Z_{s}$ denote spurious correlation factors (e.g., identity, background), and $Z_{c}$ denote causal factors induced by manipulations (e.g., forgery traces). In this SCM, the real causal path $U\!\rightarrow\!Z_{c}\!\rightarrow\!Y$ carries forgery-related causal factors, whereas $U\!\rightarrow\!Z_{s}\!\rightarrow\!Y$ constitutes a backdoor path that induces spurious correlations in training process.
@@
However, under ideal circumstance, the classifier should rely solely on forgery-related features to determine authenticity, and be invariant to interference of forgery-irrelevant factors $Z_{s}$. Since $Z_{s}$ and $Z_{c}$ are unobserved during training, we impose these desiderata at the representation level. Given an input face image $I$, we write its representation as $R=\phi(I)=h(Z_{s},Z_{c},\varepsilon)$. For brevity, the encoder $h(\cdot)$ mixes $Z_{s}$ and $Z_{c}$ into the representation $R$. We seek a representation that (i) preserves causal factor $Z_{c}$ for predicting $Y$, and (ii) blocks the backdoor path:
$\underbrace{I\!\left(Y;R\right)\ \geq\ \kappa}_{\text{causal sufficiency}}\quad\text{and}\underbrace{Y\ \perp\!\!\!\perp\ Z_{s}\ \big|\ R}_{\text{ backdoor blocking / invariance}}$ (2)
where $I(;)$ denotes mutual information, $\kappa$ is a threshold. Intuitively, the learned representation should retain enough information about the forgery-related factor $Z_{c}$ that determines $Y$, while being insensitive to spurious factors $Z_{s}$ or discard them. In the next section, we describe a transformation on $\phi(I)$ that achieves these causal desiderata.
4 Methodology
Figure 2: Pipeline of the proposed method SeLop. Based on the Structural Causal Model, we propose Low-rank Orthogonal Removal of Spurious Correlation (LROR) to mitigate influence of forgery-irrelevant factors $X_{s}$, which aims to break the backdoor path from spurious correlations to labels. See text for details.
4.1 Theoretical Analysis
To decouple spurious correlation factors from visual representations at the representation-level, we first consider separability of visual representations. According to previous section, visual representations $R=h(Z_{s},Z_{c},\varepsilon)$.
A first–order expansion yields:
$\Delta R\;\approx\;J_{s}\,\Delta z_{s}\;+\;J_{c}\,\Delta z_{c},$ (3)
where $J_{s}=\partial h/\partial z_{s}\in\mathbb{R}^{D\times m_{s}}$ and $J_{c}=\partial h/\partial z_{c}\in\mathbb{R}^{D\times m_{c}}$. We assume that $\Delta z_{s}$ and $\Delta z_{c}$ are independent and zero–mean, then the cross term vanishes and $R$ admits an additive form. The equation is expressed as:
$R\;\approx\;S+N+\xi,\quad\operatorname{Cov}\!\big(S,N\big)\approx 0,$ (4)
where $S$ approximates forgery-related representations, while $N$ denotes representations of spurious correlation. This decomposition relies on the assumption that the causal and spurious components are locally separable and approximately independent, which facilitates analytical tractability. Details are provided in supplementary material Sec.2. Moreover, (3) implies that spurious correlation variation lies in the column space of $J_{s}$:
$\operatorname{rank}\!\big(\operatorname{Cov}(N)\big)=\operatorname{rank}(J_{s}\Sigma_{s}J_{s}^{\top})\;\leq\;m_{s}\;\ll\;D.$ (5)
where $\operatorname{Cov}$ is Covariance. If spurious correlation is driven by a discrete domain variable $G\in\{1,\ldots,K\}$, the covariance admits an ANOVA decomposition $\Sigma_{R}=\mathbb{E}[\Sigma_{R\mid G}]+\sum_{k=1}^{K}\pi_{k}(\mu_{k}-\mu)(\mu_{k}-\mu)^{\top}$, whose between–domain term has rank at most K1K{-}1. Hence the spurious correlation component concentrates in a low–dimensional subspace $\mathcal{U}$ with effective rank $r\leq m_{s}+(K{-}1)\ll D$:
$R\;\approx\;S\;+\;{N}\;+\;\xi,\quad\dim(\mathcal{U})=r\ll D.$ (6)
Equations (4)–(6) summarize theoretical analysis we rely on: the causal and spurious factors are approximately separable in the visual representation, and the spurious component is low–rank compressible.
4.2 Low-rank Orthogonal Removal of Spurious Correlation
@@
According to the heuristic analysis of Sec 4.1, we propose an orthogonal low-rank projection to intervene in the visual space of CLIP. Our goal is to enable the model to adaptively and correctly remove the influence of spurious correlation features in the visual representation through training, and utilize the forgery-related features to make classification decisions.
@@
In order to retain CLIP’s rich pre-training knowledge to the maximum extent, we only apply the projection intervention in the middle and deep layers of the CLIP to estimate a spurious correlation subspace $\mathcal{U}$ (see Figure 2).
@@
where $\mathbf{Q}^{\top}\mathbf{Q}=I_{r}$, $\mathbf{Q}$ is a column-orthogonal matrix and $\mathbf{r}$ is an upper triangular matrix. Then, we decompose visual tokens into a low-rank compression subspaces $\mathcal{U}$:
$\mathcal{U}=X_{\text{s}}=X_{vis}\mathbf{Q}\mathbf{Q}^{\top}$ (8)
Due to the approximate separability and low-rank compressibility (see equation 6), forgery-irrelevant features are projected into this low-rank subspace through training. The causal factors associated with artifacts have been learned within orthogonal complement space. Specifically, its orthogonal complement space is:
$X_{\text{c}}=X_{vis}-\mathcal{U}$ (9)
Finally, we concat $X_{\text{c}}$ and $X_{[CLS]}$ to form a new token stream $X_{\text{c}},X_{[CLS]}]$ and feed it to the next transformer layer. During training, we only optimize $\mathbf{Q}$, and the final fully-connected layer while keeping the parameters of all transformer block fixed. This implementation ensures the preservation of pre-trained knowledge of CLIP while reducing the introduction of extraneous noise. Formally, our model learns $\mathbf{Q}$ to push $X_{vis}\mathbf{Q}\mathbf{Q}^{\top}$ to align with spurious variation from $Z_{s}$ and removes it, preserving causal evidence from $Z_{c}$ ($U\!\rightarrow\!Z_{c}\!\rightarrow\!Y$) and blocking the backdoor $U\!\rightarrow\!Z_{s}\!\rightarrow\!Y$.
@@
4.3 How to remove spurious correlation subspaces and why our method works

### v2

# Low-rank Orthogonal Subspace Intervention for Generalizable Face Forgery Detection

@@ Abstract
The generalization problem remains a key challenge in face forgery detection. This paper explores the reasons for the generalization failure of Vanilla CLIP: in “real vs. fake” detection, the few dominant principal components in the feature space primarily encode forgery-irrelevant information, rather than authentic forgery traces. However, this irrelevant information inevitably leads to spurious correlations, severely limiting detector performance. We define this phenomenon as “low-rank spurious bias”. To address this, we propose a low-rank representation space intervention paradigm, named the SeLop, from the perspective of causal representation learning. SeLop unifies the spurious correlation factors irrelevant to forgery into a low-rank subspace and cuts off the statistical shortcut between it and the label, thus aligning representation learning with authentic forgery traces. Specifically, we decompose spurious correlation features into a low-rank subspace through orthogonal low-rank projection, then remove this subspace from the original representation and train its orthogonal complement to capture forgery-related features. This low-rank projection removal effectively eliminates spurious correlation factors, ensuring that classification decisions are based on authentic forgery cues. With only 0.39M trainable parameters, our method achieves state-of-the-art performance across several benchmarks, demonstrating excellent robustness and generalization.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
I Introduction
@@
Figure 1: (a) Visualization of Vanilla CLIP attention mapping. We can observe that Vanilla CLIP always pays attention to forgery-irrelevant regions (e.g., identity and background) and treats them as discriminative cues. (b) PCA analysis for Vanilla CLIP’s feature space across multiple deepfake datasets. The figures show that CLIP’s 1024-dimensional high-dimensional feature space generally exhibits a low-rank manifold distribution across the datasets: the first 32 principal components alone explain over 75% of the total feature variance.
@@
To delve into this problem, we first visualize the attention of Vanilla CLIP[46] using GradCAM[50], as shown in Fig. 1 (a). Untuned CLIP tends to focus on forgery-irrelevant information (e.g., human’s identity and background), treating them as discriminative cues. This is because CLIP was pretrained on 400 million image-text pairs and did not directly learn the meaning of “forgery”, resulting in strong semantic prioritization and image semantic generalization capabilities. Furthermore, we conduct an empirical PCA study on the Vanilla CLIP representation space on five deepfake datasets, as shown in Fig. 1 (b). We randomly sample 100,000 patch features from each dataset for PCA analysis and calculate the cumulative explained variance ratio. A striking and consistent phenomenon emerges: the feature manifold exhibits a “low-rank” distribution, with a very small number of principal components explaining the vast majority of the variance (e.g., over 75% of the total variance is concentrated in the 32 principal components). This indicates that the Vanilla CLIP’s features are highly redundant. The few principal components that contribute significantly are essentially fitting natural semantic variations in real-world images, rather than real forgery traces (combined with Fig. 1 (a) and Fig. 1 (b)). However, for face forgery detection, these inherent biases often lead the model to take shortcuts and generate spurious correlations. Simultaneously, subtle forgery traces are severely suppressed and “drown” in a flat, long-tailed subspace, making standard fine-tuning difficult to capture robust, generalizable patterns. We define the above phenomenon as low-rank spurious bias.
To address this, we propose a novel approach called SeLop (Spurious correlation elimination via Low-rank orthogonal projection for face forgery detection). SeLop innovatively addresses the low-rank spurious bias phenomenon from the perspective of causal representation learning. We define forgery-irrelevant information that causes spurious correlation as spurious correlation factors, while the authentic forgery traces are causal features. SeLop’s core idea is to treat these spurious correlation factors as a whole, performing a unified subspace-level intervention in the CLIP visual representation space and cutting off the backdoor path from spurious correlation factors to groundtruth labels. Specifically, given the low-rank manifold distribution of the original CLIP feature space on deepfake data, we introduce a trainable basis matrix and construct a compact orthogonal subspace through QR decomposition. Subsequently, we explicitly remove this low-rank subspace from the original representation through orthogonal projection operations, forcing the model to mine sparse but essential causal forgery cues in the remaining orthogonal complement space. This mechanism achieves a dataset-independent unified elimination. Through end-to-end training with only 0.39M trainable parameters, the model is able to learn causal features related to forgery while preserving the rich pre-trained knowledge of CLIP to the maximum extent.
@@
    Low-rank spurious bias phenomenon in Vanilla CLIP: Through GradCAM visualization of the attention map and PCA analysis for feature space’s energy spectrum, we discover low-rank spurious bias. That is, a few significantly contributing principal components encode forgery-irrelevant information (e.g., identity and background) rather than subtle forgery traces.
@@
    We innovatively propose a simple yet effective intervention method SeLop from the perspective of causal representation learning. By applying an orthogonal low-rank intervention to the original CLIP representation space, we effectively eliminate the influence of spurious correlations unrelated to forgery traces, forcing the model to rely on true causal features for decision-making.
@@
    Extensive experiments on several standard benchmarks demonstrate the superior generalization and robustness of our method. In particular, with only 0.39M trainable parameters, our method achieves state-of-the-art performance.
The rest of this paper is organized as follows. Section II introduces the related work. In Section III, the proposed method is introduced in detail. Extensive experiment results and analysis are given in Section IV. Section V is the conclusion.
II Related Work
In this section, we introduce traditional face forgery detection methods, face forgery detection methods based on pretrained visual foundation models, and methods that incorporate Causal Representation Learning (CRL).
II-A Traditional face forgery detection
The rapid advancement of artificial intelligence technologies has spurred diversification in facial forgery techniques. Presently, improving generalizability of detection models remains a challenging issue in face forgery detection tasks. Previous work has focused on meticulously designing deep neural networks to learn “specific and fake patterns”, including biological signals, frequency anomalies, boundary artifacts and so on. LipForensics [17] trains a temporal network using fixed mouth embeddings from both real and fake data to detect fake videos based on mouth movements without overfitting to low-level, tamper-specific artifacts. DFGaze [44] uncovers subtle forgery cues by analyzing differences in facial gaze angles between real and fake videos. F3Net [45] proposes a two-stream collaborative learning framework that leverages two different but complementary frequency-aware clues to deeply mine forgery patterns. F2trans [39] combines center difference attention and wavelet transform to extract high-frequency fine-grained features, which improves detection performance effectively. Tall [58] performs fixed-position masking on consecutive frames within each frame and rearranges them into a predefined layout as thumbnails, effectively uncovering spatiotemporal clues in forged videos. FTCN-TTN [73] proposes a novel two-stage framework of temporal convolution network and temporal transformer network, aiming to explore the long-term temporal coherence.
@@ II-A Traditional face forgery detection
However, due to the limited forgery patterns in training datasets, models often overfit and fail to generalize to unknown forgery techniques. To mitigate overfitting, [27, 51] proposed a pseudo-fake face generation strategy. Training models on custom datasets improved generalization to some extent. However, expecting models to learn all forgery patterns on custom data remains impractical.
II-B Face forgery detection based on pretrained visual foundation models
@@
With the advent of large language models (LLMs), pre-trained visual foundation models (e.g., CLIP [46]) on massive data have demonstrated strong generalization across downstream tasks. Recent research has explored adapting CLIP for face forgery detection. UDD [13] designed a shuffle branch and a mix branch to break position and content biases of the CLIP. GM-DF [25] proposes a domain-aware hybrid expert modelling approach within a meta-learning framework, achieving generalized multi-scenario detection. FCG [18] has thoroughly explored spatial and temporal clues in deepfake videos. MFCLIP [71] employs textual prompts to guide the learning of forgery traces. FFD-STA [65] proposed an innovative video-level blending approach to extract spatial and temporal features. RepDFD [29] injects perturbations into input images to reprogram pre-trained VLM models. Forensic-Adapter [6] introduces an adapter specifically to learn blending boundaries, achieving excellent generalization through interaction between CLIP and the adapter. Effort [61] employs Singular Value Decomposition (SVD) to decouple the feature space of attention, preserving pre-trained knowledge while learning forgery patterns. These methods explore various ways to adjust CLIP for face forgery detection and achieve good results.
II-C Causal Representation Learning
@@
In recent years, Causal Representation Learning (CRL) [49] has attracted significant attention in the field of computer vision. Mitigating spurious correlations through causal intervention has proven a promising approach to improving cross-domain generalization [66]. Recently, some researchers have introduced the idea of CRL into face forgery detection. CADDM [12] discovered and quantified the negative impact of identity information on deepfake detection, and ultimately improved the model’s generalization ability by mitigating spurious correlation caused by identity. ID3 [67] jointly trained a purified invariant predictor and learned an aligned invariant representation, which forced the model to rely on domain-invariant features for determination, rather than being affected by forgery-irrelevant factors. UDD [13] identified position and content biases in CLIP [46], designing shuffling and mixing branches to break inherent biases and achieve great generalization. CausalCLIP [31] learned stable causal features robust to distribution shifts by modeling the generation process with a structural causal model and enforcing statistical independence. The success of these methods suggest that mitigating spurious correlations does indeed help improve generalization.
@@
Figure 2: Causal analysis. Baseline fits to spurious correlations, relying on statistical characteristics of labels and forgery-irrelevant information for classification decisions. There is a backdoor path $U\!\rightarrow\!Z_{s}\!\rightarrow\!Y$. However, our method cuts off the shortcut after successful intervention, forcing the model to truly rely on forgery traces for classification.
Figure 3: Pipeline of the proposed method SeLop. Based on the Structural Causal Model, we propose Low-rank Orthogonal Removal of Spurious Correlation (LROR) to mitigate influence of forgery-irrelevant factors $Z_{s}$, which aims to break the backdoor path from spurious correlations to labels. See text for details.
III Methodology
In this section, we first introduce our modeling approach in detail through III.A (Problem Formulation). Then, we introduce our specific method (Low-rank Orthogonal Removal of Spurious Correlation). Finally, we explain why our method effectively mitigates the impact of spurious correlations.
III-A Problem Formulation from Causal Representation Learning
@@
In this paper, we formulate the problem of face forgery detection from a novel perspective of causal representation learning. We utilize Structural Causal Model (SCM) [43] to define the relevant variables and factors, as shown in Fig. 2. The data-generating process can be described by the structural equations:
$\begin{split}Z_{s}:=g_{s}(U,\varepsilon_{s}),\quad Z_{c}:=g_{c}(U,\varepsilon_{c}),\\ R:=h(Z_{s},Z_{c},\varepsilon_{r}),\quad Y:=r(Z_{c},\varepsilon_{y}),\end{split}$ (2)
where $\varepsilon_{\cdot}$ are exogenous noises. $g_{s},g_{c},h,r$ are structural functions in the SCM. For brevity, the encoder $h(\cdot)$ mixes $Z_{s}$ and $Z_{c}$ into the representation $R$ (the image formation and the feature extractor are folded into $h$). Finally, $r(Z_{c},\varepsilon_{y})$ defines the task orientation where the ground-truth label depends on forgery cues.
Let $U$ denote unobserved latent confounders, $Z_{s}$ denote spurious correlation factors (e.g., identity, background), and $Z_{c}$ denote causal factors induced by manipulations (e.g., forgery traces). In this SCM, the real causal path $U\!\rightarrow\!Z_{c}\!\rightarrow\!Y$ carries forgery-related causal factors, whereas $U\!\rightarrow\!Z_{s}\!\rightarrow\!Y$ constitutes a backdoor path that induces spurious correlations in training process.
@@
Driven by the low-rank spurious bias phenomenon observed in Section I, we recognize that the variations of $Z_{s}$ dominate the principal components of the visual feature space, severely suppressing the subtle causal factors $Z_{c}$. However, under ideal circumstance, the classifier should rely solely on forgery-related features to determine authenticity, and be invariant to interference of forgery-irrelevant factors $Z_{s}$. Since $Z_{s}$ and $Z_{c}$ are unobserved during training, we impose these desiderata at the representation level. Given an input face image $I$, we write its representation as $R=\phi(I)=h(Z_{s},Z_{c},\varepsilon)$. Intuitively, the learned representation should retain enough information about the forgery-related factor $Z_{c}$ that determines $Y$, while being insensitive to spurious factors $Z_{s}$ or discard them. In the next section, we describe a transformation on $\phi(I)$ that achieves these desiderata.
III-B Low-rank Orthogonal Removal of Spurious Correlation
@@
To address low-rank spurious bias phenomenon and effectively block the backdoor path $U\rightarrow Z_{s}\rightarrow Y$, we propose an orthogonal low-rank projection to intervene in the visual space of CLIP. Our goal is to enable the model to adaptively and correctly remove the influence of spurious correlation features $Z_{s}$ in the visual representation through training, and utilize the forgery-related features $Z_{c}$ to make classification decisions.
@@
In order to retain CLIP’s rich pre-training knowledge to the maximum extent, we only apply the projection intervention in the middle and deep layers of the CLIP to estimate a spurious correlation subspace. (see Fig. 3).
@@
where $\mathbf{Q}^{\top}\mathbf{Q}=I_{r}$, $\mathbf{Q}$ is a column-orthogonal matrix and $\mathbf{r}$ is an upper triangular matrix. According to our PCA analysis, forgery-irrelevant signals exhibit a low-rank manifold distribution. Therefore, we can explicitly estimate the representation of spurious correlation $Z_{s}$ by decomposing the visual tokens into a low-rank compression subspace:
$Z_{\text{s}}=X_{vis}\mathbf{Q}\mathbf{Q}^{\top}$ (4)
By actively projecting the features into this low-rank subspace through training, the redundant, macroscopic semantics (e.g., identity and background) are isolated. Consequently, the causal factors $Z_{c}$ associated with authentic forgery artifacts are disentangled and preserved within the orthogonal complement space:
$Z_{\text{c}}=X_{vis}-Z_{\text{s}}$ (5)
By intervening in the visual representation, we suppress the effects of spurious correlation. Finally, we concat $Z_{\text{c}}$ and $X_{[CLS]}$ to form a new token stream $Z_{\text{c}},X_{[CLS]}]$ and feed it to the next transformer layer. See Algorithm 1 for the specific pseudocode. During training, we only optimize $\mathbf{Q}$, and the final fully-connected layer while keeping the parameters of all transformer block fixed. This implementation ensures the preservation of pre-trained knowledge of CLIP while reducing the introduction of extraneous noise. Formally, our model learns $\mathbf{Q}$ to push $X_{vis}\mathbf{Q}\mathbf{Q}^{\top}$ to align with spurious variation from $Z_{s}$ and removes it, preserving causal evidence from $Z_{c}$ ($U\!\rightarrow\!Z_{c}\!\rightarrow\!Y$) and blocking the backdoor $U\!\rightarrow\!Z_{s}\!\rightarrow\!Y$.
@@
Algorithm 1 Low-rank Orthogonal Removal of Spurious Correlation
1:Original token stream $X=[X_{vis},X_{[CLS]}]\in\mathbb{R}^{B\times(1+N_{p})\times D}$.
2:Intervened token stream
3:Output: $Z_{c},X_{[CLS]}]\in\mathbb{R}^{B\times(1+N_{p})\times D}$.
4:Learnable basis $\mathbf{B}\in\mathbb{R}^{D\times r}$, $\mathbf{Q},\mathbf{r}\leftarrow\mathrm{QR}(\mathbf{B})$ ;
5:// QR decomposition
[…]
7:// Shape:  $[B,1,D]$ and $B,N_{p},D]$
8:Define low-rank projector $P=\mathbf{Q}\mathbf{Q}^{\top},P_{\perp}=I-P$
[…]
III-C Why our method works
We model the spurious correlation subspace by a column-orthonormal basis $Q\in\mathbb{R}^{D\times r}$ with $r\ll D$, obtained by orthogonalizing a slim learnable matrix $B$ via QR decomposition to get $Q$, which induces the projector $P=QQ^{\top}$ and its orthogonal complement $P_{\perp}=I-P$.